MULTIPLE CHOICE. Choose the na t ive that best completes the statement or answer the questi 1) Write the equilibri e the equi Just give me the answers
Answer #1

1.) 2CH4 (g) + 3O2 (g) ----->2 CO(g) + 4H2O(g)

equilibrium constant = Kc = [CO]2[H2O]4/[CH4]2[O2]3

option C is correct

2) H2(g) + Br2 (g) -----> 2HBr(g)

inital = 0.682 mol 0.440 mol 0 mol

at equilibrium = (0.682 - a) (0.440- a) 2a

where a = degree of dissociation

now at equilibrium, concentration of H2 = 0.566 mol

but at equilibrium,concentration of H2 = (0.682 - a) mol

therefore, 0.682 - a = 0.566

a = 0.116 mol

concentration of Br2 at equilibrium = 0.440 - 0.116 = 0.324 mol

Option B is correct

3.)Given,

concentration of SO2 = 1.20 mol

concentration of O2 = 0.45 mol

concentration of SO3 = 1.8 mol

equilibrium constant = 5

according to question,theoretical value of Kc is

equilibrium constant = Kc = [SO3]2/[SO2]2[O2]

putting concentration values = 1.82/(1.22 x 0.45)

Kc = 5

both given and theoretical values are same,therefore system is in equilibrium

Option C is correct

4) relation between Kc and Kp is given as

Kp = Kc (RT)\bigtriangleupn

where \bigtriangleup n = (sum of molar concentration co-efficient of product side) - (sum of molar concentration coefficient of reactant side)

here \bigtriangleup n = (2+4) -(3+2) = 1

therefore Kp = Kc(RT)

therefore Kp > Kc (

option B is correct